THE DIAGONAL OF A RECTANGLE MEASURE 13 CM. THE WIDTH OF THE RECTANGLE IS 5 CM FIND THE LENGTH OF THE RECTANGLE.

THE LENGTH OF THE THE RECTANGLE IS __________CM

Answer by ewatrrr(10682)


Hi,
Applying the Pythagorean Theorem a^2 + b^2 = c^2 (diagonal being the hypotenuse)
13cm^2 - 5cm^2 = L^2
169cm^2 - 25cm^2 = L^2
144cm^2 = L^2
12cm = L
the length of the rectangle is 12cm